OpenStudy (anonymous):

Marissa has a snow cone cart that serves about 30 people each hour at the local beach. If the cone shaped snow cone cups have a height of 3 inches and a radius of 1.5 inches, what volume of shaved ice is needed each hour? Round to the nearest cubic inch. Use 3.14 for \pi .

Vocaloid (vocaloid):

total amount of ice = (number of people)*(volume of the cone)

OpenStudy (anonymous):

I got 423.9

OpenStudy (anonymous):

wait I got 45

Vocaloid (vocaloid):

volume of the cone = (1/3)(pi)r^2(h) = ?

OpenStudy (anonymous):

94.5?

Vocaloid (vocaloid):

not quite, check your math again volume of the cone = (1/3)*(height)*(pi)*(radius)*(radius) = ?

OpenStudy (anonymous):

565.2

Vocaloid (vocaloid):

not quite, let's try one more time height = 3 radius = 1.5 pi = 3.14 volume = (1/3)*(height)*(pi)*(radius)*(radius) = ?

Vocaloid (vocaloid):

* means multiply

OpenStudy (anonymous):

mabeye u should show them the fomula with everything plugged in @Vocaloid

OpenStudy (anonymous):

14.13

Vocaloid (vocaloid):

volume = (1/3)*(3)*(3.14)*(1.5)*(1.5) = ?

OpenStudy (anonymous):

5.4165

OpenStudy (anonymous):

@Aliypop what are u doing for the equation

OpenStudy (anonymous):

I'm multiplying my calculator doesn't have a fraction mode on it

OpenStudy (anonymous):

just do 1 divided by 3

OpenStudy (anonymous):

2.023333333

OpenStudy (anonymous):

no it is .3333333333333333333333333333333333333333333333333333333333333333333333333

OpenStudy (anonymous):

the answer is on the link stated by @vocaloid

OpenStudy (anonymous):

thank you :D, could you help me with 3 more questions

OpenStudy (anonymous):

if u make a new question then sure dont forget to medal